# Stepwise Regression and Pre-trained Edge for Robust Stereo Matching

### v3

@@ III-B Stepwise Regression Stereo
Figure 3: The overall architecture of the proposed SR-Stereo. Compared to iteration-based methods, SR-Stereo is specially designed in terms of the update unit and the regression objective. Specifically, a stepwise regression unit is proposed to control the range of disparity clips. Accordingly, a regression objective segmentation is proposed to design a separate regression objective for each stepwise regression unit.
Figure 3 shows the overall architecture of the proposed SR-Stereo, which is improved based on the iterative stereo method IGEV-Stereo. Intuitively, the proposed SR-Stereo reconstructs the update unit and the regression objective, and proposes a loss weight related to the regression objective scale to improve the accuracy of disparity clips.
III-B1 Feature Encoder
@@
Following IGEV-Stereo, we use MobileNetV2 [47] pre-trained on ImageNet [48] and a series of upsampling blocks to extract multi-scale features $f_{1}^{left,i}\in\mathbb{R}^{C_{i}\times H/2^{i+1}\times W/2^{i+1}}$ from the left image $I_{left}\in\mathbb{R}^{3\times H\times W}$ and $f_{1}^{right,i}\in\mathbb{R}^{C_{i}\times H/2^{i+1}\times W/2^{i+1}}$ from the right image $I_{right}\in\mathbb{R}^{3\times H\times W}$, respectively. The $i$=1,2,3,4 and the feature size $C_{i}$=48,64,192,160.
III-B2 Context Encoder
@@
We use a series of residual blocks [49] and downsampling layers to extract multi-scale context features $f_{2}^{left,j}\in\mathbb{R}^{C_{j}\times H/2^{j+1}\times W/2^{j+1}}$ ($j$=1,2,3,4 and $C_{j}$=128) from the left image. These features are inserted into the stepwise regression unit to provide global information.
III-B3 Fused Cost Volume
@@
By combining Group-wise correlation and Correlation, we construct the cost volume $G_{c}$ based on $f_{1}^{left,1}$ and $f_{1}^{right,1}$. Then, we use a lightweight 3D UNet to aggregate the cost volume and obtain the initial disparity $d_{init}\in\mathbb{R}^{1\times H/4\times W/4}$.
III-B4 Disparity Stepwise Regression
We propose a disparity stepwise regression, which obtains the disparity by predicting multiple disparity clips with the same range. It consists of two parts, including stepwise regression unit and regression objective segmentation.
@@ III-B4 Disparity Stepwise Regression
Figure 4: Architecture of the stepwise regression unit. The $m$ is the hyperparameter that controls the range of the output disparity clip.
@@
Stepwise Regression Unit. Figure 4 shows the architecture of the stepwise regression unit. For each regression, we index a set of features $f_{G}$ from the cost volume $G_{c}$ using the cumulative disparity-centered set as follows:
$f_{G}=\sum_{r=-4}^{4}concat\left\{G_{c}(d_{k}+r),G_{c}^{p}(d_{k}/2+r)\right\}$ (1)
@@
where $d_{k}$ is the cumulative disparity ($d_{0}$ = $d_{init}$) and $p$ denotes the average pooling operation. We use these features from $G_{c}$ along with the cumulative disparity $d_{k}$ to update the hidden state $h_{k}$ of ConvGRU [17]:
$h_{k+1}=ConvGRU(f_{G},d_{k},h_{k},f_{2}^{left})$ (2)
@@
Then, the updated hidden state $h_{k+1}$ and $f_{G}$ are passed into a series of convolutional layers and residual layers [49] to generate the disparity clip $\Delta d_{k+1}\in\mathbb{R}^{1\times H/4\times W/4}$ as follows:
@@
$\hfil\Delta d_{k+1}=tanh(\frac{Conv_{3\times 3}(Relu(Conv_{3\times 3}(h_{k+1})))}{m})\\ \times m\odot(1+0.5w)\hfil$ (4)
where $\sigma$ denotes the sigmoid function, $Res$ is the residual layer, and $\odot$ denotes the Hadamard product. We set the hyperparameter $m$ to constrain the size of the disparity clip and use a weight map $w$ computed from the cost-volume features $f_{G}$ to adjust the constraint magnitude. Finally, we update the current cumulative disparity and use the Spatial Upsampling [17] to generate the full-resolution cumulative disparity $d_{k+1}^{full}\in\mathbb{R}^{1\times H\times W}$ as follows:
$d_{k+1}=d_{k}+\Delta d_{k+1}$ (5)
$d_{k+1}^{full}=Upsampling(f_{1}^{left,i},I_{left},4d_{k+1})$ (6)
Regression Objective Segmentation. We specify how to obtain the ground truth disparity clip $\Delta d_{gt,k}$ and the full-resolution ground truth cumulative disparity $d_{gt,k+1}^{full}$. Specifically, the ground truth of the full-resolution cumulative disparity is constantly changing with the stepwise regression process, as shown below:
$d_{gt,k+1}^{full}=d_{k}^{full}+clip_{-6m}^{6m}(d_{gt}-d_{k}^{full})$ (7)
@@
$\Delta d_{gt,k}=clip_{-1.5m}^{1.5m}(\frac{Resize(d_{gt})}{4}-d_{k})$ (9)
where $Resize$ denotes linear interpolation downsampling.
III-B5 Disparity Clip-Balanced Weight
In SR-Stereo, the regression disparity is divided into multiple clips with the same range to mitigate the distributional discrepancies among different domains. Ideally, a regression disparity of size $L$ is split into a number of disparity clips of size $1.5m$ and a small disparity clip of size $n$ as follows:
@@ III-B5 Disparity Clip-Balanced Weight
where $k\in\mathbb{Z}$ and $n\in(0,1.5m)$. When the initial disparity is far from the ground truth disparity (i.e., $L$ is large), the distribution of disparity clips is long-tailed, which limits the accuracy of small disparity clip. In this paper, a Disparity Clip-Balanced Weight is proposed to mitigate the long-tailed distribution of disparity clips. The formula for this weight is as follows:
@@
III-B6 Loss Function
We use Smooth $L1$ loss to supervise the initial disparity and the disparity clips $d_{k}$ because it is more sensitive to large errors:
@@ III-B6 Loss Function
where $\gamma=0.9$ and $N$ is the total number of disparity clips. We use $L1$ loss to supervise full-resolution cumulative disparity:
@@
III-B7 Generalization of the stepwise regression architecture
We observe that although the core idea is different, the stepwise regression architecture is similar to existing iteration-based methods in terms of processing flow, and thus can be easily generalized based on existing methods. In Section IV-D, we also apply the proposed stepwise regression architecture to the classic iteration-based method RAFT-Stereo, and the experimental results also show that the stepwise regression architecture performs better.
III-C Domain Adaptation Based on Pre-trained Edge
We propose a Domain Adaptation Based on Pre-trained Edge (DAPE) to mitigate edge blurring for model fine-tuned with sparse ground truth. Specifically, a lightweight edge estimator is proposed to generate edge map of the target domain image, which is filtered to generate the edge map background pseudo-label. This pseudo-label, along with the sparse ground truth disparity in the target domain, are utilized as supervision to jointly fine-tune the pre-trained stereo matching model. In the following, we describe the edge estimator, the edge map background pseudo-label and the joint fine-tuning process in detail.
III-C1 Edge Estimator
For the proposed DAPE, the reliability of edge pseudo-label is vital to the performance of the fine-tuned model. Existing edge estimation methods usually take only an RGB image as input and use a series of complex 2D CNNs to estimate the contour of object by extracting high-level features. However, this approach significantly increases the computational cost and limits the generalization performance due to the large discrepancies in object classes among different domains.
Figure 5: Architecture of the edge estimator. The entire architecture consists of only one residual block and one convolution block.
In this paper, we claim that using disparity for edge estimation is a better and more robust approach because disparity estimation focuses on local geometric information, whose representations are similar across domains. Moreover, compared to RGB image, disparity contains almost no object texture information, which greatly reduces the difficulty of edge estimation. Thus, a lightweight edge estimator is proposed to achieve accurate and robust performance by introducing disparity. The architecture of the edge estimator is shown in Figure 5. First, a residual block is used to extract the edge feature in the disparity:
@@ III-C1 Edge Estimator
III-C2 Edge Map Background Pseudo-label Generation for Target Domain
@@
III-C3 Joint Fine-tuning on the Target Domain
@@
where $edge_{d_{pred}}$ is the edge map corresponding to predicted disparity $d_{pred}$. As in Section IV-F1, we use the Prewitt operator to extract the edge map of the predicted disparity:
@@ IV-A Datasets and Evaluation Metrics
KITTI 2012 [32] and KITTI 2015 [31] are datasets of real-world driving scenes. KITTI 2012 consists of 194 training pairs and 195 test pairs, while KITTI 2015 contains 200 training pairs and 200 test pairs. Both datasets provide sparse ground truth disparity (sparsest in several datasets) obtained using lidar. The disparity values range from 0 to 230. In this paper, we combine KITTI 2012 and KITTI 2015 datasets into a unified dataset referred to as KITTI. For evaluation, we adopt the 3-pixel error as the metric. To assess the efficacy of our proposed DAPE method, we allocate 80% of the KITTI training set for fine-tuning the model and reserve the remaining 20% for validation purposes.
@@ IV-B1 SR-Stereo
All ablation versions of SR-Stereo are trained on SceneFlow with a batch size of 4 for 50k steps, while the final version of SR-Stereo is trained on SceneFlow with a batch size of 8 for 200k steps. The final model and ablation experiments are conducted using a one-cycle learning rate schedule with learning rates of 0.0002 and 0.0001, respectively. We evaluate the generalization performance of the proposed method by directly testing on the 27 training pairs from ETH3D and the 15 training pairs from Middlebury 2014.
@@ IV-B2 DAPE
For the experiments related to the edge estimator, we jointly train the stereo model and the proposed edge estimator on SceneFlow with a batch size of 4 for 50k steps, using a one-cycle learning rate schedule with a learning rate of 0.0001. We use the pre-trained stereo model and edge estimator to generate edge pseudo-labels for target domains. Following existing methods [16, 22, 15], we adopte different settings of fine-tuning process for different datasets. For the KITTI, we adopt a batch size of 4 and fine-tune the model for 50k steps with an initial learning rate of 0.0001. As for the ETH3D, we use a batch size of 2 and fine-tune the model for 2,000 steps, also with an initial learning rate of 0.0001. In the case of the Middlebury 2014, we utilize a batch size of 2 and fine-tune the model for 4,000 steps, starting with an initial learning rate of 0.00002.

### v4

@@ III-B Preliminaries
Figure 3: The overall architecture of the proposed SR-Stereo. Compared to iteration-based methods, SR-Stereo is specially designed in terms of the update unit and the regression objective. Specifically, we propose a stepwise regression unit that outputs range-controlled disparity clips, rather than unconstrained residual disparities. Further, we design separate regression objectives for each stepwise regression unit, instead of simply using the disparity error.
Essentially, the core of the iteration-based methods is the regression of the residual disparity $\Delta d_{k}$ to the disparity error $d_{gt}-d_{k-1}^{*}$. In contrast, in our SR-Stereo, the disparity error is split into multiple range-controlled segments for regression, thus mitigating the domain discrepancies, as shown in Figure 4. In this paper, we reconstruct the update unit and regression objective in existing iteration-based methods to implement SR-Stereo. To highlight the differences, we refer to the output of the new update unit as a disparity clip rather than residual disparity.
Figure 4: Visualization of the regression objectives of SR-Stereo and iteration-based methods. The iteration-based methods regress disparity error by predicting residual disparity $\Delta d_{k}$, while SR-Stereo splits the disparity error into multiple segments and regresses them by predicting multiple disparity clips.
III-C Stepwise Regression Stereo
In this section, we describe the key components that implement SR-Stereo and show how they can be plugged into iteration-based methods. Using IGEV-Stereo [10] as the basic model, the overall architecture is shown in Figure 3. Intuitively, we retain the original feature extraction and cost-volume construction, but reconstruct the update unit and its regression objective. In addition, we propose a loss weight for the new update unit that is related to the regression objective scale.
III-C1 Feature extraction and Cost volume Construction
We do not change the feature extraction and cost-volume construction of IGEV-Stereo [10], but still describe them in this subsection.
@@ III-C1 Feature extraction and Cost volume Construction
Feature Encoder. We use MobileNetV2 [29] pre-trained on ImageNet [30] and a series of upsampling blocks to extract multi-scale features $f_{1}^{left,i}\in\mathbb{R}^{C_{i}\times H/2^{i+1}\times W/2^{i+1}}$ from the left image $I_{left}\in\mathbb{R}^{3\times H\times W}$ and $f_{1}^{right,i}\in\mathbb{R}^{C_{i}\times H/2^{i+1}\times W/2^{i+1}}$ from the right image $I_{right}\in\mathbb{R}^{3\times H\times W}$, respectively. The $i=1,2,3,4$ and the feature size $C_{i}=48,64,192,160$.
@@
Context Encoder. We use a series of residual blocks [31] and downsampling layers to extract multi-scale context features $f_{2}^{left,j}\in\mathbb{R}^{C_{j}\times H/2^{j+1}\times W/2^{j+1}}$ ($j$=1,2,3,4 and $C_{j}$=128) from the left image. These features are inserted into the stepwise regression unit to provide global information.
@@
Cost Volume. By combining Geometry Encoding Volume (GEV) and All-pairs Correlations (APC), we construct the cost volume $G_{c}$ based on $f_{1}^{left,1}$ and $f_{1}^{right,1}$. Then, we use a lightweight 3D UNet to aggregate the GEV and obtain the initial disparity $d_{init}\in\mathbb{R}^{1\times H/4\times W/4}$.
III-C2 Stepwise Regression Architecture
In this subsection, we detail the two key components that implement SR-Stereo: the stepwise regression unit and the regression objective segment.
@@ III-C2 Stepwise Regression Architecture
Figure 5: Architecture of the stepwise regression unit. The $m$ is the hyperparameter that controls the range of the output disparity clip. The $\Delta d_{k,ori}$ is the residual disparity output of the original update unit, and $Res$ is the residual layer.
@@
Stepwise Regression Unit. Figure 5 shows the architecture of the stepwise regression unit, which highlights the improvements over the original update unit. For each regression, we index a set of features $f_{G}$ from the cost volume $G_{c}$ using the current disparity-centered set as follows:
$f_{G}=\sum_{r=-4}^{4}concat\left\{G_{c}(d_{k-1}+r),G_{c}^{p}(d_{k-1}/2+r)\right\}$ (5)
@@
where $d_{k-1}$ is the current disparity ($d_{0}$ = $d_{init}$) and $p$ denotes the average pooling operation. We use these features from $G_{c}$ along with the current disparity $d_{k-1}$ to update the hidden state $h_{k-1}$ of ConvGRU [10]:
$h_{k}=ConvGRU(f_{G},d_{k-1},h_{k-1},f_{2}^{left})$ (6)
@@
where $h_{k}$ is the updated hidden state. Then, the $h_{k}$ and $f_{G}$ are passed into a series of convolutional layers and residual layers [31] to generate the disparity clip $\Delta d_{k}\in\mathbb{R}^{1\times H/4\times W/4}$ as follows:
$\hfil\Delta d_{k}=tanh(\frac{\Delta d_{k,ori}}{m})\times m\odot(1+0.5w)\hfil$ (7)
in which
@@
$\Delta d_{k,ori}=Conv_{3\times 3}(Relu(Conv_{3\times 3}(h_{k})))$ (9)
where $\sigma$ denotes the sigmoid function, $Res$ is the residual layer, $\odot$ denotes the Hadamard product, and $\Delta d_{k,ori}\in\mathbb{R}^{1\times H/4\times W/4}$ is the residual disparity output of the original update unit. We set the hyperparameter $m$ to constrain the approximate range of the disparity clips, and use the weight map $w$ to adaptively adjust the constraint magnitude for different regions. The weight map $w$ is computed from the cost-volume feature.
Finally, we update the current disparity and use the Spatial Upsampling [10] to generate the full-resolution disparity $d_{k}^{full}\in\mathbb{R}^{1\times H\times W}$ as follows:
$d_{k}=d_{k-1}+\Delta d_{k}$ (10)
$d_{k}^{full}=Upsampling(f_{1}^{left,i},I_{left},4d_{k})$ (11)
Regression Objective Segment. We specify how to obtain the ground truth disparity clip $\Delta d_{gt,k}$ and the full-resolution ground truth disparity $d_{gt,k}^{full}$. Specifically, the ground truth of the full-resolution disparity is constantly changing with the stepwise regression process, as shown below:
$d_{gt,k}^{full}=d_{k-1}^{full}+clip_{-6m}^{6m}(d_{gt}-d_{k-1}^{full})$ (12)
@@
$\Delta d_{gt,k}=clip_{-1.5m}^{1.5m}(\frac{Resize(d_{gt})}{4}-d_{k-1})$ (14)
where $Resize$ denotes linear interpolation downsampling. We point out that supervision of the disparity clips is not necessary, but slightly improves the model’s performance (see Section IV-C for experimental results).
Figure 6: Long-tailed distribution of regression objectives for disparity clips. SR-Stereo regresses the disparity error by predicting multiple disparity clips, but the regression objectives of disparity clips are long-tailed.
III-C3 Disparity Clip-Balanced Weight
As mentioned earlier, SR-Stereo mitigates the domain discrepancies by predicting multiple disparity clips to regress the disparity error. However, this approach leads to a new problem: the regression objective for the disparity clips exhibits a long-tailed distribution, as shown in Figure 6. Ideally, a disparity error of size $L$ can be regressed by multiple clips of size $1.5m$ and a small clip of size $n$ as follows:
@@ III-C3 Disparity Clip-Balanced Weight
where $k\in\mathbb{Z}$ and $n\in(0,1.5m)$. Obviously, the distribution of regression objectives is long-tailed (especially when the disparity error is large), which limits the accuracy of small disparity clips. In this paper, a Disparity Clip-Balanced Weight is proposed to mitigate the long-tailed distribution. The formula for this weight is as follows:
@@
Figure 7: The overall framework of the proposed DAPE. First, a robust stereo model SR-Stereo and a lightweight edge estimator are pre-trained on a large synthetic dataset with dense ground truth. Then, we use the pre-trained SR-Stereo and edge estimator to generate the edge map of target domain, where the background pixels (i.e., non-edge region pixels) are used as edge pseudo-labels. Finally, we jointly fine-tune the pre-trained SR-Stereo using the edge pseudo-labels and sparse ground truth disparity.
III-C4 Loss Function
We use Smooth $L1$ loss to supervise the initial disparity and the disparity clips $\Delta d_{k}$:
@@ III-C4 Loss Function
where $\gamma=0.9$ and $N$ is the total number of disparity clips. We use $L1$ loss to supervise full-resolution disparities:
@@
III-D Domain Adaptation based on Pre-trained Edge
We propose a Domain Adaptation based on Pre-trained Edge (DAPE) to mitigate edge blurring for model fine-tuned with sparse ground truth. As shown in Figure 7, the proposed DAPE consists of three steps:
  1. 1.
    Firstly, a robust stereo model and a lightweight edge estimator are pre-trained on a large synthetic dataset (i.e., the source domain) with dense ground truth. Specifically, the input to the edge estimator is the disparity predicted by the stereo model along with the RGB image.
  2. 2.
    Then, the pre-trained stereo model and the edge estimator are used to directly infer on the target domain and generate corresponding edge maps. To mitigate the negative impact of disparity errors in reflective regions, we recommend using only the relatively dense background pixels (i.e., non-edge region pixels) from the edge maps as pseudo-labels (following UCFNet [14], these pixels are defined as pseudo-labels).
  3. 3.
    After generating the edge pseudo-labels of the target domain, we use it and the sparse ground truth disparity to jointly fine-tune the pre-trained stereo model.
In the following, we describe the edge estimator, the edge map background pseudo-labels and the joint fine-tuning process in detail.
III-D1 Edge Estimator
For the proposed DAPE, the reliability of edge pseudo-label is vital to the performance of the fine-tuned model. Existing edge estimation methods usually take only an RGB image as input and use a series of complex 2D CNNs to estimate the contour of object by extracting high-level features. However, this approach significantly increases the computational cost and limits the generalization performance due to the substantial differences in object categories and distributions across different domains.
Figure 8: Architecture of the edge estimator. The entire network consists of only two residual blocks and one convolution block.
In this paper, we claim that using disparity for edge estimation is a better and more robust approach because disparity estimation focuses on local geometric information, whose representations are similar among different domains. Moreover, compared to RGB image, disparity contains almost no object texture information, which greatly reduces the difficulty of edge estimation. Thus, a lightweight edge estimator is proposed to achieve accurate and robust edge prediction by introducing disparity. The architecture of the edge estimator is shown in Figure 8. First, a residual block is used to extract the edge feature in the disparity:
@@ III-D1 Edge Estimator
III-D2 Edge Map Background Pseudo-label Generation for Target Domain
@@
III-D3 Joint Fine-tuning on the Target Domain
@@
where $edge_{d_{pred}}$ is the edge map corresponding to predicted disparity $d_{pred}$. As in Section IV-E1, we use the Prewitt operator to extract the edge map of the predicted disparity:
@@ IV-A Datasets and Evaluation Metrics
KITTI 2012 [19] and KITTI 2015 [18] are datasets of real-world driving scenes. KITTI 2012 consists of 194 training pairs and 195 test pairs, while KITTI 2015 contains 200 training pairs and 200 test pairs. Both datasets provide sparse ground truth disparity (sparsest in several datasets) obtained using lidar. The disparity values range from 0 to 230. In Section IV-E2, we merge KITTI 2012 and KITTI 2015, referred to as KITTI. For evaluation, we adopt the 3-pixel error as the metric. To assess the efficacy of our proposed DAPE method, we allocate 80% of the KITTI training set for fine-tuning the model and reserve the remaining 20% for validation.
TABLE I: Ablation study of SR-Stereo. The baseline (i.e. *) is IGEV-Stereo. The $num_{GRU}$ denotes the usage times of ConvGRU, while the $num_{SRU}$ denotes the usage times of the proposed stepwise regression unit. All methods run 15 disparity updates during inference ($num_{GRU}+num_{SRU}=15$). The final configuration of $m$ is underlined. Bold: Best. Blue: Second.
Index Variations SceneFlow Middlebury-H ETH3D Params.(M)
$num_{GRU}$ $num_{SRU}$ $m$ $Loss_{\Delta d}$ EPE(px) $>3px$ $>2px(\%)$ $>1px(\%)$
(a)* 15 0 - - 0.72 3.65 8.44 4.49 12.60
(b) 0 15 1 - 0.71 3.61 8.42 4.10 12.77
(c) 0 15 2 - 0.69 3.51 7.89 4.47 12.77
(d) 0 15 3 - 0.70 3.55 7.66 4.44 12.77
(e) 0 15 4 - 0.69 3.54 7.31 5.14 12.77
(f) 0 15 2 $\surd$ 0.70 3.49 6.78 4.05 12.77
(g) 0 15 3 $\surd$ 0.69 3.49 7.44 4.15 12.77
(h) 4 11 2 $\surd$ 0.69 3.49 7.50 4.61 12.77
(i) 4 11 3 $\surd$ 0.70 3.56 7.80 4.61 12.77
(j) 8 7 2 $\surd$ 0.69 3.49 7.77 4.10 12.77
(k) 8 7 3 $\surd$ 0.70 3.52 7.91 4.79 12.77
@@ IV-B1 SR-Stereo
In the experimental section, following Selective-Stereo [11] and MoCha-Stereo [33], we implement the proposed SR-Stereo on the basis of IGEV-Stereo for a fair comparison. All ablation versions of SR-Stereo are trained on SceneFlow with a batch size of 4 for 50k steps, while the final version of SR-Stereo (represented by SR-IGEV) is trained on SceneFlow with a batch size of 8 for 200k steps. The final model and ablation experiments are conducted using a one-cycle learning rate schedule with learning rates of 0.0002 and 0.0001, respectively. Following some generalized stereo methods [39, 38, 40], we directly evaluate cross-domain performance on the training sets of KITTI, Middlebury, and ETH3D.
@@ IV-B2 DAPE
For the experiments related to the edge estimator, we jointly train the stereo model and edge estimator on SceneFlow with a batch size of 4 for 50k steps, using a one-cycle learning rate schedule with a learning rate of 0.0001. Then, We use the pre-trained stereo model and edge estimator to generate edge pseudo-labels for target domains. Following existing methods [9, 13, 8], we adopt different settings of fine-tuning process for different datasets. For the KITTI, we adopt a batch size of 4 and fine-tune the model for 50k steps with an initial learning rate of 0.0001. As for the ETH3D, we use a batch size of 2 and fine-tune the model for 2,000 steps, also with an initial learning rate of 0.0001. In the case of the Middlebury 2014, we utilize a batch size of 2 and fine-tune the model for 4,000 steps, starting with an initial learning rate of 0.00002.